OpenStudy (callisto):

Co-geometry question #1

OpenStudy (callisto):

Sorry, my brain is not functioning well at the moment. Here's an extremely easy question

OpenStudy (callisto):

AB = 4/3 and AC =1 and the coordinates of B and C are given

OpenStudy (experimentx):

seems like you can get x and y solving the distance equations http://www.wolframalpha.com/input/?i=solve+%28x%2B5%2F3%29%5E2+%2B+y%5E2+%3D+%284%2F3%29%5E2%3B+x%5E2%2By%5E2+%3D+1 slope will be quite easy.

OpenStudy (callisto):

Yup, this is one of the methods, but there is another way to do it, and my question is from there. It's using tan

OpenStudy (experimentx):

must be some other special elegant method.

OpenStudy (callisto):

Indeed, a very cute solution :)

OpenStudy (callisto):

slope = tan (angle ABC) = 1/ (4/3) = 3/4

OpenStudy (experimentx):

Now problem is how do you find that angle??

OpenStudy (callisto):

I understand how it comes. But not sure when does it work. When should I use tangent ratio?

OpenStudy (callisto):

the problem is I don't know when to use tan ratio to find slope and when I can't

OpenStudy (experimentx):

Oo ... did i hear something like 3/4 is the answer??

OpenStudy (callisto):

No, but you saw :P

OpenStudy (experimentx):

Lol ... I hadn't realized it yet. I'm quite thick headed, I never listen to anyone. I think you can take tan ratio as slope when your base is parallel to x-axis.

OpenStudy (callisto):

what do you mean by 'base' ?

OpenStudy (experimentx):

 If \(\theta\) is the reference angle, then b is base, if this b is parallel to x-axis then we can take tan theta as slope.